OpenStudy (anonymous):

Find the gradient of each side of this quadrilateral (0,0) (12,5) (0,10) (-6,8) I will draw it

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

gradient is another word for slope to find the slope of each side, use the slope formula m = (y2-y1)/(x2-x1) So let's find the slope of the line going through (0,0) and (12,5) m = (y2-y1)/(x2-x1) m = (5-0)/(12-0) m = 5/12 So the slope of the line that passes through the two points (0,0) and (12,5) is 5/12 --------------------------------------------------------------------------------------------------------------- Now let's find the slope of the line going through (12,5) and (0,10) m = (y2-y1)/(x2-x1) m = (10-5)/(0-12) m = 5/(-12) m = -5/12 So the slope of the line that passes through the two points (12,5) and (0,10) is -5/12 --------------------------------------------------------------------------------------------------------------- I'll let you do the rest.

OpenStudy (anonymous):

OOOh yeah I remember doing that now thank you :) But how also does one go about finding the length of a side?

jimthompson5910 (jim_thompson5910):

Use the distance formula so for example, say we want to find the distance from (12,5) to (0,10) d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((0-12)^2+(10-5)^2) d = sqrt((-12)^2+(5)^2) d = sqrt(144+25) d = sqrt(169) d = 13 So exact distance between the two points (12,5) and (0,10) is 13 units.